Find the covariance and correlation coefficient for the following sets of data. Select the answers equal to or closest to your results. X: 50 44 47 40 54 Y: 10 13 95 7 Cov What does each measure tell you? Check all that apply. The covariance tells you that there is a weak or nonexistent linear relationship between X and Y The covariance and correlation coefficient tell you that there is a positive linear relationship between X and Y. The correlation coefficient tells you that there is a negative linear relationship between X and Y □ The correlation coefficient tells you that there is a strong linear relationship between X and Y □ The correlation coefficient tells you that there is a weak or nonexistent linear relationship between X and Y. What dont these measures tell you? Check all that apply. If you observe a larger value of X, you are likely to see a larger value of Y associated with it. If you change the value of X, you will make the value of Y become smaller. A larger value of X causes the value of Y to be larger. If you change the value of X, you wl make the value of Y become larger.
Answer #1

Answer: Cov= 1.25

r= 0.0765

Correlation measures tells that there is a weak or non-existence of linear relationship between X and Y as it is very less or less than 0.1.

In second case all of the options are correct, as there is a weak or non-existence of relationships so we can not say anything about X and Y.

Calculation Sum(X) = 50 + 44 +47 + 40 + 54 = 235 Mean = 47 Sum(Y) 10 13+ 9+5+ 7 -44 YMean - 8.8 Covariance(X,Y) = SUM (Xi-X,mean)*(y Ymean)/(samplesize = (50-47)*(10-8.8)+ (44-47)*(13-8.8)+ (47-47)*(9-8.8)+ (40-47)*(5-8.8)+ (54-47)*(7-8.8))/4 = 1.25